OpenStudy (anonymous):

Estimate the area under the graph f(x)=6/x, from x=1 to x=7 using 6 approximate rectangles and right endpoints.

OpenStudy (anonymous):

since the length of the path from 1 to 7 is 6, this is going to be easy each interval is of length one, and your right hand endpoints are 2,3,4,5,6,7 the base for each rectangle is 1, and multiplying by 1 is a lot like doing nothing so your actual job is to compute \[6\left(\frac{1}{2}+\frac{1}{3}+\frac{1}{4}+\frac{1}{5}+\frac{1}{6}+\frac{1}{7}\right)\] i would use a calculator

OpenStudy (anonymous):

Thank you so much!!! I was having so much trouble with this one!

OpenStudy (anonymous):

Awesome! It was correct!